Mini split #1 Lennox Tech found system really low on R410a, suspects 4-5lbs. Found a leak at the fitting on the airhandler flare fitting. Tech spoke with Ben P. About the findings and approved the repairs on the Lennox mini split. Tech tightened down flare fitting and leak tested with soap bubbles to verify that the leaks have stopped. Tech changed system with 4lbs of R410a to acquire proper refrigerant levels and delta T as well. System is back cooling. Approval over the phone.
Tech found the heat pump condenser #6 with black oil Marks thru the condenser coil, compressor plug disconnected, service disconnect off, and one of terminals is blown out at the compressor. Office will be informed of the findings.
Performed cma full service on 12 split systems and 2 mini split systems. Performed inspection on oven hood exhaust and make up air. Mini split system 1: Nearby landscaping growing in to unit. Recommend trimming of bush for air flow. System not cooling properly. Recommend demand tech to further diagnose. Mini split system 2: provided estimate for replacement of batteries in remote/thermostat. System 1: Heat pump exposed to excessive lint from nearby laundry vent. System 2: Air handler auxiliary drain pan holding water. Recommend installing ez trap for future service of drain line. Provided estimate for installation $138.12. Removed ~9 gal water from auxiliary pan. Flushed drain line with compressed nitrogen. Informed maintenance to monitor for future problems. Outdoor fan motor capacitor tested weak. Provided estimate for replacement $200.50. System 3: Heat pump condenser coil shows signs of rupture. Leak detector shows signs of refrigerant leak. Recommend replacement of condensin
Tech found 2 shorted heat strip contactors on 2 different systems, system #8 and #2. #2 had a tripped transformer at the airhandler due to the heat strip contactor shorted. Tech spoke with Ben P. About findings and approved replacing 2 heat strip contactors on both systems, #2,#8. Tested system operation afterwards. Both systems that were worked on are operating properly at this time.